It takes 15ml of 0.5 M NaOH solution to completely neutralize 5ml of H2SO4 solution. What is the concentration of the H2SO4 solution?
Answer #1

For a neutralisation reaction,we use the formula c1v1=c2v2 ---------(i)

Here,v1= 15 ml NaOH=(15/1000)L NaOH ; c1= 0.5 M

v2= 5 ml H2SO4= (5/1000)L H2SO4 ; c2= ?

From formula (i),

(15*0.5)/1000 = (5*c2)/1000

   c2=1.5 M

Concentration of H2SO4 is 1.5 M.

We were unable to transcribe this image

We were unable to transcribe this image